question 1 -

A voltaic cell is constructed from a standard Ag+|Ag half cell (E°red = 0.799V) and a standard Hg2+|Hg half cell (E°red = 0.855V).

The anode reaction (including states) is:


The cathode reaction(including states) is:


The spontaneous cell reaction is:

The cell voltage is ....

question 2

A voltaic cell is constructed from a standard Cr3+|Cr half cell (E°red = -0.740V) and a standard Fe3+|Fe2+ half cell (E°red = 0.771V).

The anode reaction (including states) is:


The cathode reaction(including states) is:


The spontaneous cell reaction is:

The cell voltage is ....

question 3 -

A voltaic cell is constructed from a standard Al3+|Al half cell (E°red = -1.660V) and a standard Sn2+|Sn half cell (E°red = -0.140V).

The anode reaction (including states) is:


The cathode reaction(including states) is:


The spontaneous cell reaction is:

The cell voltage is ....
Answer #1

1. Ans - Giver half cell an Aat/A ... Esed = 0.7990 149 27/4 ... Esed = 0.8550 from the standard electrode potential groen ,STAO cathode and Get 1 Gr. So, fest is ar is oxidiced at reduced at anode . Anode reaction - Cr(s) o taft e f Cathode reactio:30A9 STAGI PAGE : DATE: I 1 2Al(s) +35m?+ (aq) → Quo137 (aq) + 35m (5) We have E cell = E cathode - Eanode = -0.140-1 -2.660